# Distribution Aligned Semantics Adaption for Lifelong Person Re-Identification

### v1

@@ 3.3 Overall Pipeline
The framework is shown in Fig. 2. At each training step, we use the basic identity classification loss $\mathcal{L}_{id}$ [33] for supervision. After training the model, we can obtain the domain-specific $BN^{(t)}$ and $SA^{(t)}$ as introduced above. Both of them are lightweight and stored for taking advantage of the learned semantics at the inference stage. Note that the extra parameters introduced per domain are not significant compared with other competitors as shown in Fig. 4. During testing, according to the camera information of the person image, corresponding $BN^{(t)}$ and $SA^{(t)}$ are used. With the help of lightweight $BN^{(t)}$ and $SA^{(t)}$, the model can efficiently evolve and adapt to new application domains, making it effective and practical for new deployment and wide usage in real-world scenarios.
@@ 4.1 Experimental Setup
Datasets. We evaluate our proposed method following the widely-used LReID setting [14, 12, 13]. Specifically, we investigate its effectiveness in two different training orders on various person Re-ID datasets. “Order 1”: Market-1501 [35] $\rightarrow$ DukeMTMC-reID [36] $\rightarrow$ CUHK-SYSU [37] $\rightarrow$ MSMT17 [38]. “Order 2”: VIPeR [39] $\rightarrow$ Market-1501 $\rightarrow$ CUHK-SYSU $\rightarrow$ MSMT17. Note that DukeMTMC-reID is only used for academic use and fair comparison without identifying or showing pedestrian images. CUHK-SYSU is modified from the original for person search and rearranged following [14]. We split VIPeR into training and testing sets following [13]. For other datasets, we follow their original training and evaluation protocols. To investigate its generalization capability, we also evaluate it on the unseen datasets, including CUHK01 [40], CUHK02 [41], GRID [42], SenseReID [43], PRID [44].
We also evaluate the influence of using different datasets for pre-training, including ImageNet [24] for the image classification task, as well as LUPerson [8] and LUPerson-NL [9] for person Re-ID. LUPerson is a large-scale unlabeled dataset of 4M images of over 200K identities. LUPerson-NL is derived from LUPerson by applying a simple online multi-object tracking system on the raw videos of LUPerson. LUPerson-NL consists of 10M images with about 430K identities collected from 21K scenes. Unless otherwise specified, we use the LUPerson-NL pre-trained weights for our proposed method, which are effective in providing robust and general human semantics as shared knowledge for lifelong evolution.
Evaluation metrics. Mean Average Precision (mAP) and Rank-1 accuracy (R-1) are adopted for performance evaluation in seen domains after each training step, and report average accuracies by averaging mAP and R-1 in all seen domains, denoted as $\overline{s}_{\rm mAP}$ and $\overline{s}_{\rm R-1}$. We also evaluate the performance in unseen domains after the last training step.
@@ 4.2 Implementation Details
We adopt ResNet-50 [10] as the backbone model following previous works [11, 14, 12, 13], and the input images are resized to $256\times 128$ with random horizontal flipping, padding, random cropping, and random erasing [45] for data augmentation. The batch size is set to 128, with 2 samples per pedestrian. Adam optimizer  [46] with weight decay of $5\times 10^{-4}$ is adopted, with the warmup strategy that linearly increases the learning rate from $3.5\times 10^{-5}$ to $3.5\times 10^{-4}$ in the first 10 epochs. The learning rate is decreased by a factor of 10 at the 30-$th$ epoch for the first dataset and at the 10-$th$ epoch for other datasets. Each dataset is trained for 80 epochs. For the SA module, the kernel size is set to 5$\times$ 5 with stride 1, and the bias term is removed for the sake of the number of parameters and its limited effect.
@@ 4.3 Comparison with State-of-the-Arts
Results in seen domains. We compare our proposed method with advanced lifelong learning methods for classification and recognition [15, 17, 18, 19, 20, 21, 22] and those specially designed for LReID [11, 14, 13, 12]. Results of training Orders 1 and 2 are shown in Tabs. 1 and 2, respectively. Results of comparison methods for LReID are reproduced in the same experimental environment according to their released code. Fine-Tune denotes fine-tuning on different datasets sequentially. Joint-Train denotes combining all datasets for joint training.
As shown in Tab. 1, due to the discrepancies across domains, although rehearsal-based methods show great performance in the classification task [19, 20], it contributes to inferior performance gain when being applied to LReID. For the methods designed for LReID, AKA [11] struggles to maintain a balance between forward knowledge learning and retention. Other advanced methods [12, 13, 14] find it effective to draw support from exemplars to promote performance following the knowledge distillation paradigm. Without relying on impractical exemplars, our proposed DASA outperforms them significantly. As shown in Fig. 3, DASA also achieves better average accuracies during the lifelong evolution process on the four datasets. The results demonstrate its effectiveness in adapting to the application domain and learning domain-specific knowledge. With the catastrophic forgetting problem being effectively avoided, it is effective and applicable in real-world LReID scenarios.
@@
As shown in Tab. 2, our proposed DASA also achieves state-of-the-art results when training in Order 2. On the small-scale VIPeR [39] dataset with only 632 person images for training, DASA still shows great superiority. By comparing the results of Orders 1 and 2, we have the following observations. (1) DASA achieves better results on Market-1501 when using Order 2 compared with Order 1 where Market-1501 is used as the first domain. It is owing to the design that the previously learned $BN^{(t-1)}$ and $SA^{(t-1)}$ are used as the initial to train $BN^{(t)}$ and $SA^{(t)}$ effectively, achieving the forward transfer of knowledge from the learned domain to the next. (2) The training order has a great impact on the state-of-the-art competitors. For example, PTKP [12] achieves slightly better results than KRC [13] in Order 1, but much worse results in Order 2. Another example is that on Market-1501, PTKP shows much better results in Order 1 compared with Order 2 (surpassing about 20% on mAP and R-1). (3) KRC performs better than PTKP on the small-scale VIPeR. In contrast, affected by the exemplars from VIPeR, PTKP achieves bad incremental accuracies on the following datasets. These results show the great sensitivity of advanced LReID methods to the chosen training datasets. (4) The catastrophic forgetting problem is not well addressed by other advanced methods. For instance, KRC performs worse on Market-1501 in Order 1 since more new domains participate in the lifelong learning process than those in Order 2. In contrast, DASA is robust to different training orders and datasets.
@@
Comparison of storage consumption. Storage consumption is also crucial in lifelong scenarios, otherwise, we can save all previous models to prevent catastrophic forgetting. We compare the storage consumption for different LReID methods after the last training step (Order 1) in Fig. 4 and Tab. 3. Our proposed method shows significant superiority over advanced competitors. All competitors [12, 13, 14] have to store the classifier for knowledge distillation to relieve forgetting. As introduced in Sec. 3.1, as the process of lifelong learning, the dimension of the identity category increases dramatically. The classifier, implemented as one fully connected layer, would consume more and more, whose size is almost seven-tenths of the size of the model after the last training step as shown in Tab. 3. What’s worse, an old model having the same classifier with a large dimension is used for knowledge distillation at each training step, further exacerbating storage consumption.
@@
Results in unseen domains. Our proposed DASA shows great anti-forgetting capability and outstanding performance in application domains. We also investigate its generalization capability in unseen domains. We follow the competitors to use the model at the last training step which accumulates rich knowledge from all seen domains. As shown in Tabs. 5 and 5, the generalization ability of DASA outperforms that of all other advanced methods as well as Fine-Tune and Joint-Train on all datasets and in both training orders. The results indicate that robust and general human semantic knowledge is useful for generalization and DASA shows great effectiveness in generalizing acquired knowledge.
Table 4: Comparison with state-of-the-art methods in unseen domains after training in Order 1. The results (Rank-1) are given by the trained model after the last training step.
@@
Table 5: Comparison with state-of-the-art methods in unseen domains after training in Order 2. The results (mAP) are given by the trained model after the last training step.
@@ 4.4 Ablation Studies
For a fairer comparison and to validate our designs in LReID, we conduct all ablation experiments using the LUPerson-NL [9] pre-trained weights in Tab. 6. Besides state-of-the-art methods, we also compare the results of (1) Fine-Tune: fine-tuning the pre-trained model on different datasets sequentially; (2) DA: tuning BN layers to achieve distribution alignment at each training step; (3) SA: freezing the pre-trained model and only adding and tuning our proposed SA modules; (4) our proposed DASA; (5) Joint-Train: combining all datasets for joint training. Note that for SA, since BNNeck [33] is not trained during pre-training, it is also trained at each training step. In Fig. 6, we further investigate the influence of various pre-training choices on different methods, that is adopting different pre-trained weights trained on the ImageNet [24], LUPerson [8], and LUPerson-NL [9] datasets for model initialization, respectively. We also explore the choices of kernel sizes for the SA module, the positions to use SA, and the effectiveness of SA in adapting semantics via attention map visualization. Without loss of generality, ablation results are conducted in Order 1 by default.
Table 6: Ablation Studies of our method. All methods use LUPerson-NL pre-trained weights. The results are reported after the last training step in Order 1.
@@
Effectiveness of the SA module. The results in Tab. 6 show that SA also significantly beats all advanced competitors [11, 13, 12] using the same pre-trained weights. It demonstrates that just using the pre-trained weights from large-scale person datasets for model initialization cannot effectively promote LReID. The knowledge distillation paradigm limits their performance upper bound. However, our proposed SA is effective in leveraging human semantics learned from pre-training to promote LReID.
Necessity of adopting both distribution alignment and semantics adaption. As shown in Tab. 6, combining both DA and SA achieves the best results. On the one hand, although the pre-trained model can acquire human semantic knowledge to represent pedestrians, the data distribution discrepancy between pre-training and application would result in ineffective use of the learned semantics for specific applications. On the other hand, with the help of the SA module for aggregating semantic representations after distribution alignment, on the challenging DukeMTMC-reID and MSMT17 datasets, DASA brings 6.4% and 9.5% mAP improvement over DA, respectively. Note that DASA achieves competitive results with Joint-Train, and even better mAP and R-1 on the CUHK-SYSU dataset. We guess it is because the image styles and data distributions of CUHK-SYSU differ greatly from others. When trained jointly, the model has to make a compromise between datasets. On the contrary, without being influenced by distribution discrepancy, DASA leverages acquired knowledge of human semantics with the SA modules to represent pedestrians effectively.
@@
Where to use the SA module. We also investigate the influence of applying SA in different layers. Considering grid search on each layer takes a great effort, we explore adopting it in different layers of ResNet-50 [10] in Tab. 7. (1) By comparing Methods 2$\sim$4, we find that the SA module brings a better effect in deep layers. It is consistent with our motivation since the deep layers are more specialized in semantics learning while the shallow layers are better at capturing curves and image patterns. (2) Methods 2$\sim$4 all outperform Method 1, showing the effectiveness of SA in adapting acquired knowledge for better pedestrian representation. (3) Method 5 shows competitive performance with “Ours”, and it costs even less resource. Considering the better performance, we simply apply the lightweight SA module after each frozen Conv layer. It would be also effective to adopt it only in the deep layers of the model for more efficient deployment and application.
@@
Influence of different pre-training data. DASA shows great superiority over other advanced LReID methods using the same pre-trained weights as shown in Tab. 6. We also investigate the influence of different pre-training choices on it. According to the results of Fig. 6, we observe that (1) Methods based on ImageNet pre-training generally achieve inferior results. Since ImageNet includes all kinds of objects, pre-training on it brings less useful human semantics for LReID. (2) DA using LUPerson pre-trained weights outperforms that using ImageNet pre-trained weights, showing that pre-training on the person dataset is helpful to acquire shared knowledge for LReID. However, it still performs worse than that using LUPerson-NL pre-trained weights. The results demonstrate that when the person dataset used for pre-training is not large-scale, the model cannot acquire robust and general human semantics effectively. (3) SA outperforms DA when using ImageNet pre-trained weights. The results show that on the one hand, the model does not acquire useful knowledge for LReID, since it is still suboptimal despite aligning the data distribution via DA. On the other hand, the SA module can aggregate acquired knowledge for object classification to promote LReID to a certain extent, demonstrating its effectiveness. (4) DA achieves competitive or even better results with SA when using LUPerson or LUPerson-NL pre-trained weights, showing the nonnegligible impact of the data distribution gap. (5) Aligning distribution or adopting SA can facilitate LReID in different pre-training choices. Combining them shows great superiority over Fine-Tune, and good pre-trained weights can expand the upper bound of DASA. It is appealing to promote LReID with our proposed paradigm following the trend of large-scale pre-training.
@@
Visualization of attention maps. To intuitively understand how the SA module achieves semantics adaption from pre-training to the application domain, we visualize the attention maps in Fig. 7. We observe that (1) Since the Market-1501 dataset is relatively easy with greater inter-class differences, the model is capable of taking advantage of learned general human semantics to effectively focus on the human body. As shown in the left of Fig. 7 (b), the model can distinguish pedestrians based on their different body semantics. When adopting SA, as shown in the left of Fig. 7 (c), the model concentrates more on discriminative body regions and does not rely on the face to distinguish pedestrians, which improves its robustness since the face is usually not considered the dominant information in Re-ID. Paying excessive attention to it can even be misleading due to low resolution, view changes, or occlusions. (2) MSMT17 is much more challenging with great intra-class variance and minor inter-class differences. Without semantics adaption, the model is susceptible to the background or not robust person clues, e.g., small logos or corners of clothes as shown in the right of Fig. 7 (b). However, the SA module helps to concentrate on discriminative regions of the human body, effectively promoting Re-ID. Overall, our proposed SA module can handle different domains with great discrepancy, showing great ability and prospect to evolve the model for new deployment and wide application.

### v2

@@ 3.3 Overall Pipeline
The framework is shown in Fig. 2. At each training step, we use the basic identity classification loss $\mathcal{L}_{id}$ [33] for supervision. After training the model, we can obtain the domain-specific $BN^{(t)}$ and $SA^{(t)}$ as introduced above. Both of them are lightweight and stored for taking advantage of the learned semantics at the inference stage. Note that the extra parameters introduced per domain are not significant compared with other competitors as shown in Fig. 4. During testing, according to the camera information of the person image, corresponding $BN^{(t)}$ and $SA^{(t)}$ are used. With the help of $BN^{(t)}$ and $SA^{(t)}$, the model can efficiently evolve and adapt to new application domains, making it effective and practical for new deployment and wide usage in real-world scenarios.
@@ 4.1 Experimental Setup
Datasets. We evaluate our proposed method following the widely-used LReID setting [16, 14, 15]. Specifically, we investigate its effectiveness in two different training orders on various person Re-ID datasets. Order 1: Market-1501 [44] $\rightarrow$ DukeMTMC-reID [45] $\rightarrow$ CUHK-SYSU [46] $\rightarrow$ MSMT17 [47]. Order 2: VIPeR [48] $\rightarrow$ Market-1501 $\rightarrow$ CUHK-SYSU $\rightarrow$ MSMT17. Note that DukeMTMC-reID is only used for academic use and fair comparison without identifying or showing pedestrian images. CUHK-SYSU is modified from the original for person search and rearranged following [16]. We split VIPeR into training and testing sets following [15]. For other datasets, we follow their original training and evaluation protocols. To investigate its generalization capability, we also evaluate it on the unseen datasets, including CUHK01 [49], CUHK02 [50], GRID [51], SenseReID [52], PRID [53].
We also evaluate the influence of using different datasets for pre-training, including ImageNet [25] for image classification, LUPerson [10] and LUPerson-NL [11] for person Re-ID. LUPerson is a large-scale unlabeled dataset of 4M images of over 200K identities. LUPerson-NL is derived from LUPerson by applying an online multi-object tracking system on the raw videos of LUPerson. LUPerson-NL consists of 10M images with about 430K identities collected from 21K scenes. Unless otherwise specified, we use the LUPerson-NL pre-trained weights for our method, which can provide robust and general human semantics as shared knowledge for lifelong evolution.
Evaluation metrics. Mean Average Precision (mAP) and Rank-1 accuracy (R-1) are adopted for performance evaluation on seen domains after each training step. We also report average accuracies by averaging mAP and R-1 on all seen domains (denoted as $\overline{s}_{\rm mAP}$ and $\overline{s}_{\rm R-1}$), as well as performance on unseen domains after the last training step.
@@ 4.2 Implementation Details
We adopt ResNet-50 [12] as the backbone model following previous works [13, 16, 14, 15], and the input images are resized to $256\times 128$ with random horizontal flipping, padding, random cropping, and random erasing [54] for data augmentation. The batch size is set to 128, with 2 samples per pedestrian. Adam optimizer [55] with weight decay of $5\times 10^{-4}$ is adopted, with the warmup strategy that linearly increases the learning rate from $3.5\times 10^{-6}$ to $3.5\times 10^{-4}$ in the first 10 epochs. The learning rate is decreased by a factor of 10 at the 30th epoch for the first dataset and at the 10th epoch for other datasets. Each dataset is trained for 80 epochs. For the SA module, the kernel size is set to 5 $\times$ 5 with stride 1, and the bias term is removed for the sake of the number of parameters and its limited effect.
@@ 4.3 Comparison with State-of-the-Arts
Results on seen domains. We compare our proposed method with advanced lifelong learning methods for classification and recognition [17, 19, 20, 21, 22, 23, 24], and those specially designed for LReID [13, 16, 15, 14]. The results of training Orders 1 and 2 are shown in Tabs. 1 and 2, respectively. Comparison methods for LReID are reproduced in the same experimental environment according to their released code and training settings. Fine-Tune denotes fine-tuning on different datasets sequentially. Joint-Train denotes combining all datasets for joint training. As shown in Tab. 1, due to the discrepancies across domains, although rehearsal-based methods show great performance in the classification task [21, 22], it contributes to inferior performance gain when being applied to LReID. For the methods designed for LReID, AKA [13] struggles to maintain a balance between forward knowledge learning and retention. Other advanced methods [14, 15, 16] draw support from exemplars following the knowledge distillation paradigm. Without relying on impractical exemplars, DASA still outperforms them significantly.
The previously acquired knowledge will not be forgotten in our LReID paradigm. However, previous methods forget more and more with the increasing of subsequent domains. Previous studies [13, 14] have discussed the forgetting phenomenon of the model on seen datasets. Since our LReID paradigm avoids forgetting, we compare the average accuracies at different training steps, which more comprehensively evaluates the capabilities in knowledge preservation and knowledge updating, both of which are essential for LReID. As shown in Fig. 3, DASA achieves better average accuracies during the lifelong evolution process on the four datasets. The results show its effectiveness in adapting to the application domains and acquiring knowledge.
@@
As shown in Tab. 2, our proposed DASA also achieves state-of-the-art results when training in Order 2. On the small-scale VIPeR [48] dataset with only 632 person images for training, DASA still shows great superiority. By comparing the results of Orders 1 and 2, we have the following observations. (1) The training order has a great impact on the state-of-the-art competitors. For example, PTKP [14] achieves slightly better results than KRC [15] in Order 1, but much worse results in Order 2. Another example is that on Market-1501, PTKP shows much better results in Order 1 compared with Order 2 (surpassing about 20% on mAP and R-1). (2) KRC performs better than PTKP on the small-scale VIPeR. In contrast, affected by the exemplars from VIPeR, PTKP achieves bad incremental accuracies on the following datasets. These results show the great sensitivity of advanced LReID methods to the chosen training datasets. (3) The catastrophic forgetting problem is not well addressed by other advanced methods. For instance, KRC performs worse on Market-1501 in Order 1 since more new domains participate in the lifelong learning process than those in Order 2. In contrast, DASA is robust to different training orders and datasets.
@@
Comparison of storage consumption. Storage consumption is also crucial in lifelong scenarios, otherwise, we can save all previous models to prevent catastrophic forgetting. We compare the storage consumption for different LReID methods after the last training step (Order 1) in Fig. 4 and Tab. 3. All competitors [14, 15, 16] have to store the classifier for knowledge distillation to relieve forgetting. As introduced in Sec. 3.1, as the process of lifelong learning, the dimension of the identity category increases dramatically. The classifier, implemented as one fully connected layer, would consume more and more, whose size is almost seven-tenths of the size of the model after the last training step as shown in Tab. 3. What’s worse, an old model having the same large classifier is used for knowledge distillation, further exacerbating storage consumption.
@@
Results on unseen domains. Our proposed DASA shows great anti-forgetting capability and outstanding performance in application domains. We also investigate its generalization capability on unseen domains. We follow the competitors to use the model at the last training step which accumulates rich knowledge from all seen domains. As shown in Tabs. 5 and 5, the generalization ability of DASA outperforms that of all other advanced methods as well as Fine-Tune and Joint-Train on all datasets in both training orders. The results indicate that robust and general human semantic knowledge is useful for generalization and DASA shows great effectiveness in generalizing acquired knowledge.
Table 4: Comparison with state-of-the-art methods on unseen domains after training in Order 1. The results (Rank-1) are given by the trained model after the last training step.
@@
Table 5: Comparison with state-of-the-art methods on unseen domains after training in Order 2. The results (mAP) are given by the trained model after the last training step.
@@ 4.4 Ablation Studies
For a fairer comparison and to validate our designs, we conduct all ablation experiments using the LUPerson-NL [11] pre-trained weights in Tab. 6. Without loss of generality, ablation experiments are conducted in Order 1 by default. Besides state-of-the-art methods, we also compare (1) Fine-Tune: fine-tuning the pre-trained model on different datasets sequentially; (2) DA: tuning BN layers to achieve distribution alignment at each training step; (3) SA: freezing the pre-trained model and only adding and tuning our proposed SA modules; (4) our proposed DASA; (5) Joint-Train: combining all datasets for joint training. Note that for SA, since BNNeck [33] is not trained during pre-training, it is also trained at each training step. In Fig. 6, we further investigate the influence of various pre-training choices on different methods. We also explore the choices of kernel sizes for the SA module, the positions to use SA, and the effectiveness of SA in adapting semantics via attention map visualization.
Table 6: Ablation Studies of our method. All methods use LUPerson-NL pre-trained weights. The results are reported after the last training step in Order 1. “*” denotes adding the SA module to the original ResNet-50.
@@
Effectiveness of the SA module. The results in Tab. 6 show that SA also significantly beats all advanced competitors [13, 15, 14] using the same pre-trained weights. It demonstrates that just using the pre-trained weights from large-scale person datasets for model initialization cannot effectively promote LReID. The knowledge distillation paradigm limits their performance upper bound. However, our proposed SA is effective in leveraging human semantics learned from pre-training to promote LReID. We also try to add the SA module to the ResNet-50 backbone. For Fine-Tune, the previously learned semantics are destroyed and tuned to fit the current domain, so incorporating the SA module may lead to inflexible knowledge updating and severe knowledge forgetting, resulting in performance degradation. In contrast, SA shows its great effectiveness in our proposed LReID paradigm. Additionally, since Joint-Train can acquire all data from all application domains, the SA module can also help improve semantics with few parameters. With proper design and usage, SA can bring great benefits.
Necessity of adopting both distribution alignment and semantics adaption. As shown in Tab. 6, combining both DA and SA achieves the best results. On the one hand, although the pre-trained model can acquire human semantic knowledge to represent pedestrians, the data distribution discrepancy between pre-training and application would result in ineffective use of the learned semantics for specific applications. On the other hand, with the help of the SA module for aggregating semantic representations after distribution alignment, on the challenging DukeMTMC-reID and MSMT17 datasets, DASA brings 6.4% and 9.5% mAP improvement over DA, respectively. Note that DASA achieves competitive results with Joint-Train, and even better mAP and R-1 on the CUHK-SYSU dataset. We guess it is because the image styles and data distributions of CUHK-SYSU differ greatly from others. When trained jointly, the model has to make a compromise between datasets. On the contrary, without being influenced by distribution discrepancy, DASA leverages acquired knowledge of human semantics with the SA module to represent pedestrians effectively.
@@
Where to use the SA module. Considering grid search on each layer takes a great effort, we explore adopting SA in different layers of ResNet-50 [12] in Tab. 7. (1) By comparing Methods 2$\sim$4, we find that the SA module brings a better effect in deep layers. It is consistent with our motivation since the deep layers are more specialized in semantics while the shallow layers are better at capturing curves and image patterns. (2) Methods 2$\sim$4 all outperform Method 1, showing the effectiveness of SA in adapting acquired knowledge for better pedestrian representations. (3) Method 5 shows competitive performance with Ours with even less resource cost. Considering the better performance, we apply the lightweight SA module after each frozen Conv layer. It would be also effective to adopt it only in the deep layers for more efficient deployment and application.
@@
Influence of different pre-training data. According to the results of Fig. 6, we observe that (1) Methods based on ImageNet pre-training generally achieve inferior results. Since ImageNet includes all kinds of objects, pre-training on it brings less useful human semantics for LReID. (2) DA using LUPerson pre-trained weights outperforms that using ImageNet pre-trained weights, showing that pre-training on the person dataset provides useful knowledge for LReID. However, it still performs worse than that using LUPerson-NL pre-trained weights. The results demonstrate that when the person dataset used for pre-training is not large-scale, the model cannot acquire robust and general human semantics effectively. (3) SA outperforms DA when using ImageNet pre-trained weights. On the one hand, the model does not acquire useful knowledge for LReID, since it is still suboptimal despite aligning the data distribution via DA. On the other hand, the SA module can aggregate acquired knowledge for object classification to promote LReID to a certain extent, demonstrating its effectiveness. (4) DA achieves competitive or even better results with SA when using LUPerson or LUPerson-NL pre-trained weights, showing the nonnegligible impact of the data distribution gap. (5) Aligning distribution or adopting SA can facilitate LReID in different pre-training choices. Combining them shows great superiority over Fine-Tune, and good pre-trained weights can expand the upper bound of DASA. It is appealing to promote LReID with our paradigm following the trend of large-scale pre-training.
@@
Visualization of attention maps. To intuitively understand how the SA module achieves semantics adaption from pre-training to the application domain, we visualize the attention maps in Fig. 7. We observe that (1) Since Market-1501 is relatively easy with greater inter-class differences, the model can leverage learned general human semantics to focus on the human body. As shown in the left of (b), the model can distinguish pedestrians based on their different body semantics. When adopting SA, as shown in the left of (c), the model concentrates more on discriminative body regions and does not rely on the face, which improves its robustness since the face is usually not considered the dominant information in Re-ID. Paying excessive attention to it can be misleading due to low resolution, view changes, or occlusions. (2) MSMT17 is much more challenging with greater variances. Without semantics adaption, the model is susceptible to background or fragile person clues, e.g., small logos, or corners of clothes, as shown in the right of (b). However, the SA module helps to concentrate on discriminative regions of the human body, effectively promoting Re-ID. Overall, our proposed SA module can handle different domains with great discrepancy, showing great ability and prospect to evolve the model for new deployment and wide application.